OpenStudy (anonymous):

Find the vertex of the parabola formed by the following quadratic function: y = 7x2 + 21x – 1 Answer (-2, -15) (2, 69) (1, 27) (-3/2, -67/4)

OpenStudy (anonymous):

please someone help me

OpenStudy (anonymous):

is it the last one

OpenStudy (anonymous):

no problem u can put many questions

OpenStudy (anonymous):

Shakir can you stop it now ??

OpenStudy (anonymous):

@best.shakir everytime you are giving links..

OpenStudy (anonymous):

ok sorry

OpenStudy (anonymous):

@waterineyes u can explain

OpenStudy (anonymous):

i dont know to explain so that i sent links

OpenStudy (anonymous):

Yes if you give me a chance before posting the direct answer links..

OpenStudy (anonymous):

Here, we have an quadratic equation: \[y = 7x^2 + 21x - 1\] Here: a = 7, b = 21 and c = -1: Vertex is given by: \[Vertex(h,k) = (\frac{-b}{2a}, \frac{-D}{4a})\]

OpenStudy (anonymous):

\[h = \frac{-21}{14} \implies h = \frac{-3}{2}\]

OpenStudy (anonymous):

Now find Discriminant: \(D = b^2 - 4ac = 469\) \[k = \frac{-469}{28} = \frac{-67}{4}\]

OpenStudy (anonymous):

\[(h,k) = (\frac{-3}{2}, \frac{-67}{4})\]

OpenStudy (anonymous):

Getting @best.shakir

OpenStudy (anonymous):

yep......